The following genotypes were observed in a population:

Genotype: JJ (40), Jj (45), jj (50)

a. Calculate the observed genotypic and allelic frequencies for this population.

b. Calculate the expected numbers of individuals for each genotype if this population were in Hardy- Weinberg equilibrium.

c. Using a chi-square test, determine whether the population is in Hardy-Weinberg equilibrium.
Answer #1

p2=JJ=40/135=0.30

2pq=Jj=45/135=0.33

q2=jj=50/135=0.37

Frequency of J

p=p2+1/2 2pq=0.3+1/2x0.33=0.465

p+q=1

0.465+q=1

q=1-0.465

=0.535

JJ(p2)=(0.465)2=0.216

Jj(2pq)=2x0.465x0.535=0.497

jj(q2)=(0.535)2=0.286

b)Expected no of individuals

JJ=0.216X135=29

Jj=0.497x135=67

jj=0.286x135=39

c)CHI SQUARE(X2)

X2=(o-e)2/e